IDENTIFICATION OF THE [10.3] ^{2}Φ_{7/2} - A ^{2}Δ_{5/2} TRANSITION OF NICKEL CHLORIDE

\underline{LEAH O’BRIEN} and KRYSTAL GIBBS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared system of NiCl with a bandhead at 10,168 cm^{−1}has been recorded and identified as the [10.3] ^{2}Φ_{7/2} - A ^{2}Δ_{5/2} transition. Molecular constants for the new [10.3] ^{2}Φ_{7/2} state will be presented. A comparison of the known energy levels for nickel halides will be discussed.